a) Derive the MUx and MUy b) Derive the MRSy
Answer #1

a) utility function is

U(x,y) = \alphax\delta /\delta +\betay\delta /\delta

Therefore, MUx = dU/dx = \alpha\deltax\delta-1 /\delta= \alphax\delta-1

MUy = dU/dy = \beta\deltay\delta-1 / \delta= \betay\delta-1

b) MRSxy = MUx /MUy = \alphax\delta-1 /\betay\delta-1